OpenStudy (anonymous):

What types of functions combine to create an explicit formula for arithmetic sequences?

OpenStudy (freckles):

what does your question mean ?

OpenStudy (freckles):

\[a,a+d,a+2d,a+3d,... ,a+(n-1)d,... \text{ where the } nth \text{ term is given by } \\ a_n=a+(n-1)d\]

OpenStudy (freckles):

by the way the above is an arithemetic sequence because the common difference is d

OpenStudy (freckles):

that is term - previous term will always give you d

OpenStudy (anonymous):

functions that allow the calculation of any term in the sequence ??|

OpenStudy (freckles):

you can use the above formula to find a term a_k if you know d and the first term

OpenStudy (freckles):

for example say your first term is 5 and your common difference is 3 and you want to find a_(55) \[a_n=5+(n-1)3 \\ \text{ well } a_{55}=5+(55-1)3=5+(54)3=5+162=167\]